OpenStudy (anonymous):

Recurrence Relations: Can you use generating functions to solve the recurrence relation T(n) = 2T(n-1) + n ?

OpenStudy (anonymous):

\[T_1=2(0)+1=1\] \[T_2=2(1)+2,T_3=2(4)+3,T_4=2(9)+4...\] \[T_1=1,T_2=4,T_3=9,T_4=22...\]

OpenStudy (anonymous):

i am nt sure what recuurance relation means